OpenStudy (anonymous):

Express the complex number in trigonometric form. 2 - 2i

OpenStudy (jdoe0001):

so, you don't have the formulas for converting complex to trigs?

OpenStudy (anonymous):

@jdoe0001 no

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

@jdoe0001 can you help?

OpenStudy (jdoe0001):

\[a+bi = r(cos(\theta)+i\ sin(\theta))\] now how to find out what "r" is? \[r=\sqrt{a^2+b^2}\] now, how to find out what the angle/theta is? \[tan(\theta)=\frac{b}{a}\]

OpenStudy (jdoe0001):

what signs "a" and "b" have will tell you what quadrant they're in, and thus you use THAT angle with THAT tangent

OpenStudy (jdoe0001):

in this particular case, 2-2i, a=(2), b=(-2), so b[y coordinate] is negative, and a[x coordinate] is positive, that means, where X is positive and Y is negative :), that quadrant, that angle with that tangent

OpenStudy (anonymous):

@jdoe0001 that's all the quadrants

OpenStudy (jdoe0001):

mmm, no all quadrants have such neg/pos combination, the signs actually alternate some from quadrant to quadrant

OpenStudy (jdoe0001):

take a peek at your Unitary Circle, if you have any

OpenStudy (anonymous):

@jdoe0001 quadrant 3 has it

OpenStudy (jdoe0001):

so, that's your quadrant :)

OpenStudy (anonymous):

@jdoe0001 what now

OpenStudy (jdoe0001):

well, you need to find "r" and the tangent :), using the formulas

OpenStudy (jdoe0001):

once found, just combine them in the \[r(cos(\theta)+i\ \ sin(\theta))\] form, and that's the converted version

OpenStudy (jdoe0001):

confused still?

OpenStudy (anonymous):

@jdoe001 no I got it ... 2√(2)(cos((5π)/(4))+isin((5π)/(4))) ?

OpenStudy (jdoe0001):

well, looks actually fine, more or less, but tan(5pi/4) = 1, no -1, your tangent is -2/2 = -1

OpenStudy (jdoe0001):

and also 5pi/4 is on the 3rd quadrant

OpenStudy (jdoe0001):

ohh, I did say it was so... well, my bad :|, is not the 3rd anyway :)

OpenStudy (jdoe0001):

bear in mind that \[tan(\theta)=\frac{-2}{2}=-1\]

OpenStudy (anonymous):

@jdoe0001 now what

OpenStudy (jdoe0001):

check your unit circle, see where Y is negative and X is positive, is not the 3rd quadrant though

OpenStudy (jdoe0001):

thus is not 5pi/4, the "r" is fine though

OpenStudy (anonymous):

@jdoe0001 quadrant 4

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

@jdoe0001 7pi/4

OpenStudy (jdoe0001):

yep, that's what I got